need help solving this problem
(t+3)(t+8)=t(t-5)+168

Answer by mananth(16946)

(t+3)(t+8)=t(t-5)+168
t(t+8)+3(t+8)=t^2-5t+168
t^2+8t+3t+24=t^2-5t+168
t^2+11t+24=t^2-5t+168
16t=168-24
16t=144
/16
t=144/16
t=9